# LV-Eval: A Balanced Long-Context Benchmark with

### v1

@@ Acknowledgments
Table 6: Ablation results on hotpotwikiqa-mixup for confusing facts insertion (CFI) and keyword and phrase replacement (KPR).
@@
Table 7: Ablation results on multifieldqa-en-mixup for confusing facts insertion (CFI) and keyword and phrase replacement (KPR).
@@
Table 8: Ablation results on multifieldqa-zh-mixup for confusing facts insertion (CFI) and keyword and phrase replacement (KPR).
@@
Table 9: Ablation results on factrecall-en and factrecall-zh for confusing facts insertion (CFI) and keyword and phrase replacement (KPR).
Appendix A Samples in LV-Eval
@@
Figure A12: Prompts used for GPT-4 to generate confusing facts.
Appendix B Detailed evaluation results
Dataset Len. Ch. Bl. Yi. Lo. Ll.32 Qw. Vi. Ll.4 GPT3. GPT4.
dureader -mixup $16k$ 23.99 19.40 2.87 13.44 11.82 12.00 9.67 7.21 8.01 19.14
$32k$ 25.21 19.74 2.98 11.57 10.65 12.80 7.65 5.42 5.26 13.64
$64k$ 22.01 14.44 2.88 9.23 8.58 10.48 6.62 5.59 4.26 12.66
$128k$ 17.94 10.95 2.36 9.51 9.34 8.15 6.25 4.78 3.30 8.19
$256k$ 8.72 8.51 3.06 7.96 7.48 8.65 5.70 4.45 3.50 6.71
loogle-CR -mixup $16k$ 14.41 9.01 8.25 11.25 3.11 5.48 5.00 3.69 10.04 12.68
$32k$ 14.10 7.36 8.83 11.17 2.82 3.30 4.25 3.29 8.39 10.40
$64k$ 9.92 3.81 4.73 9.31 2.01 3.82 3.76 3.13 5.58 6.48
$128k$ 6.95 2.40 4.05 6.19 2.46 1.14 1.99 2.19 3.08 2.83
$256k$ 5.46 2.60 3.23 5.03 2.16 1.94 1.28 0.81 3.37 3.91
loogle-MR -mixup $16k$ 15.83 4.90 6.94 10.53 3.12 4.93 5.17 3.37 12.95 12.24
$32k$ 11.62 3.14 7.67 9.51 2.61 2.95 3.83 2.20 7.03 7.83
$64k$ 7.00 1.68 2.69 3.04 1.44 2.37 0.96 2.05 6.23 6.26
$128k$ 7.24 2.46 3.44 4.05 1.47 1.80 0.55 1.04 2.13 2.30
$256k$ 3.82 2.19 1.32 3.01 0.95 1.46 1.06 0.33 1.00 0.90
hotpotwikiqa -mixup $16k$ 16.98 19.31 23.55 11.57 3.54 2.78 2.63 3.99 11.96 13.51
$32k$ 14.76 14.07 18.94 10.71 2.31 1.89 2.19 1.30 6.66 10.62
$64k$ 9.02 9.63 9.94 4.77 2.20 2.27 2.05 1.84 3.27 6.67
$128k$ 8.31 7.71 7.66 5.49 1.86 2.37 1.04 0.81 4.23 4.13
$256k$ 6.68 5.40 2.01 2.37 1.62 1.82 1.85 0.75 3.30 2.36
lic-mixup $16k$ 24.15 20.75 5.37 15.45 10.55 6.05 8.34 2.48 7.65 13.69
$32k$ 22.27 12.68 6.25 10.02 8.87 6.07 4.81 0.99 4.42 5.86
$64k$ 14.33 5.00 7.19 4.54 3.41 4.21 2.52 0.48 3.07 3.23
$128k$ 8.30 3.03 5.56 2.47 1.85 4.34 2.36 0.42 0.87 1.90
$256k$ 6.07 4.11 6.24 2.14 1.66 3.19 1.99 0.73 1.65 1.70
Table A1: Overall results of multi-hop QA tasks in LV-Eval. The abbreviations “Ch.”, “Bl.”, “Yi.”, “Lo.”, “Ll.32”, “Qw.”, “Vi.”, “Ll.4”, “GPT3.”, and “GPT4.” stand for ChatGLM3-6B-32k, BlueLM-7B-32k-Chat, Yi-6B-200k, LongChat-7B-32k-v1.5, Llama2-7B-32k-Instruct, Qwen-7B-8k-Chat, Vicuna-7B-16k-v1.5, Llama2-7B-Chat-hf, GPT-3.5-16k, and GPT-4-8k, respectively.
Dataset Len. Ch. Bl. Yi. Lo. Ll.32 Qw. Vi. Ll.4 GPT3. GPT4.
loogle-SD -mixup $16k$ 41.82 34.34 39.56 27.42 13.94 10.54 8.79 6.75 31.67 27.01
$32k$ 30.31 15.10 36.48 18.21 10.58 4.70 4.90 2.61 18.56 14.01
$64k$ 19.07 4.95 31.71 12.09 5.53 2.40 3.07 2.58 10.41 8.00
$128k$ 11.34 5.32 25.71 9.11 4.80 3.25 4.24 2.04 5.74 5.14
$256k$ 8.92 5.41 12.37 5.97 3.30 3.02 2.39 1.24 3.56 1.48
cmrc-mixup $16k$ 51.21 45.89 1.05 20.99 13.86 11.13 11.75 3.85 12.19 14.67
$32k$ 46.34 19.53 0.35 10.77 7.31 5.32 6.55 1.08 6.00 3.33
$64k$ 20.71 10.66 0.84 8.97 4.10 4.68 5.04 1.72 3.57 5.31
$128k$ 14.16 7.06 1.58 3.77 2.95 3.81 2.75 1.64 2.73 3.81
$256k$ 8.38 4.51 2.54 3.75 2.40 4.09 4.13 1.54 1.32 2.68
multifieldqa -en-mixup $16k$ 25.40 11.82 10.01 12.02 8.03 7.66 6.29 8.81 18.78 19.00
$32k$ 12.78 6.34 9.24 7.58 4.96 3.61 4.32 5.55 11.59 12.69
$64k$ 12.32 8.38 8.83 7.84 4.12 5.23 2.79 1.58 7.38 8.30
$128k$ 9.89 5.29 5.98 3.11 3.90 3.64 2.51 2.54 7.95 7.25
$256k$ 4.24 4.78 4.69 4.22 2.13 2.44 1.28 1.49 3.21 3.54
multifieldqa -zh-mixup $16k$ 32.38 22.05 2.85 9.81 4.55 8.82 5.82 4.72 18.94 17.61
$32k$ 24.48 17.64 0.75 8.82 3.93 5.68 4.45 1.21 12.21 11.18
$64k$ 20.97 7.36 1.89 3.23 1.45 3.01 2.03 0.68 6.29 4.99
$128k$ 10.08 5.90 2.11 3.54 1.74 2.84 0.88 0.24 2.94 1.76
$256k$ 7.05 4.48 1.58 3.92 1.15 2.52 1.26 0.56 2.15 0.92
factrecall-en $16k$ 91.50 58.5 24.88 9.22 75.20 1.77 0 1.08 8.25 23.4
$32k$ 89.00 32.17 23.09 14.33 56.00 1.12 0 0.46 3.27 11.84
$64k$ 46.00 15.50 24.96 8.31 33.00 0.71 0 0.31 1.80 5.21
$128k$ 24.00 9.00 22.04 7.86 17.85 0.18 0.25 0.23 0.60 4.03
$256k$ 12.50 5.00 16.44 6.00 8.40 0.22 0.20 0.15 0.45 1.79
factrecall-zh $16k$ 0 19.00 25.73 7.20 2.55 15.75 0 0 14.51 28.03
$32k$ 2.00 37.00 16.86 5.00 0.74 6.00 0 0 6.70 15.24
$64k$ 12.50 20.00 12.41 3.50 0.53 3.50 0 0 2.49 8.08
$128k$ 9.00 12.50 10.13 3.70 0.49 1.50 0 0 1.72 3.58
$256k$ 7.00 5.50 4.62 2.00 0.29 0.50 0 0 0.98 2.00
Table A2: Overall results of single-hop QA tasks in LV-Eval. The abbreviations “Ch.”, “Bl.”, “Yi.”, “Lo.”, “Ll.32”, “Qw.”, “Vi.”, “Ll.4”, “GPT3.”, and “GPT4.” stand for ChatGLM3-6B-32k, BlueLM-7B-32k-Chat, Yi-6B-200k, LongChat-7B-32k-v1.5, Llama2-7B-32k-Instruct, Qwen-7B-8k-Chat, Vicuna-7B-16k-v1.5, Llama2-7B-Chat-hf, GPT-3.5-16k, and GPT-4-8k, respectively.

### v2

@@ 4.2 Ablation Study of LV-Eval Techniques
5 Limitations
LV-Eval does not encompass other task types such as summarization. Additionally, due to the high cost, we do not test some of the most recent LLMs, such as GPT-4-128k, GPT-4o, and so on. As we release all the test data, one can intentionally overfit the benchmark by training on the test data to get a high score. In this case, training on LV-Eval datasets with KPR might lead to mistakes in common-sense knowledge, resulting in a very unreliable evaluation.
A worth-noting issue about KPR is that as KPR modifies the fact, there exist some cases where the model identifies factual errors and then insists on providing a common-sense response. We view these cases as an issue with the model’s instruction-following ability, as our prompt explicitly states that the answer should be solely based on the context instead of its existing knowledge.
Our CFI technique relies on manual revision of the confusing facts to ensure the benchmark quality. A possible way to alleviate the human efforts when applying the CFI technique to future corpus might be leveraging stronger LLMs to substitute manual revision efforts.
Acknowledgments
We thank the insightful discussion with Kai Chen and Songyang Zhang from Shanghai Artificial Intelligence Laboratory. Thanks to all of the annotators who contribute to the project. We thank Hanling Zhang for helping run the experiment for Moonshot.
Appendix
[…]
Figure A6: Keyword-recall-based two-stage metric calculation. (a) The vanilla F1 score (red) is inflated high. With the keyword recall-based metric, the final score is set to zero due to the low recall of answer keywords. (b) The vanilla F1 score (red) is inflated high due to irrelevant words. By filtering of blacklisted words, the final score is better calibrated.
Figure A7: Prompts used for GPT-4 to generate confusing facts.
Appendix A Detailed Construction of QA Pairs
Multi-hop QA.
In a multi-hop QA task, the reasoning to derive the answer needs to gather multiple pieces of information from various locations in the context. We construct four multi-hop QA datasets: dureader-mixup, loogle-CR-mixup, loogle-MR-mixup, and hotpotwikiqa-mixup.
Single-hop QA.
In a single-hop QA task, only a single evidence in the context is needed to derive the answer. We construct seven single-hop QA datasets: lic-mixup, loogle-SD-mixup, cmrc-mixup, multifieldqa-en-mixup, multifieldqa-zh-mixup, factrecall-en, and factrecall-zh.
[…]
  • factrecall-en and factrecall-zh are two synthetic datasets designed to assess the LLMs’ ability to identify a small piece of evidence (“fact”) located at various locations within a lengthy context. As shown in Figure A10 A11, we write one English fact-question-answer pair for factrecall-en and one Chinese fact-question-answer pair for factrecall-zh. distracting documents are sourced from PG-19 dataset (English) and the book of Journey to the West (Chinese) to create five contexts of different length levels. For each context, we generate 200 documents by inserting the fact at 200 evenly spaced positions within the context.
Appendix B Annotation Details
  • Annotators: We hire a total of five annotators, including three master students involved in LLM research and two master students majored in linguistics. They are hired to work on-site as full-time annotators.
  • Annotation time: On average, annotators worked 8 hours per day. (1) For CFI, six datasets comprising a total of 557 instances of confusing facts were reviewed, with the verification process completed by two annotators over a period of three days. This task was assigned to master’s students in linguistics to ensure semantic consistency. (2) For KPR, six datasets containing 1,924 pairs were processed. Replacing key words and phrases in the Chinese datasets involved five individuals over three days, whereas the English datasets required two individuals over two days. (3) For answer keyword annotation, nine datasets comprising 955 instances of answer keywords were annotated by two individuals within a single day.
  • Annotation guidelines: We provided guidelines and examples to the five annotators and conducted a trial annotation in accordance with these guidelines to ensure they understand the guidelines. The guidelines are summarized as follows. For CFI, two types of cases require manual modification: (a) The confusing facts generated by GPT-4 fail to produce interference; (b) The confusing facts generated by GPT-4 conflict with the original facts.
    For KPR, the guidelines are: (a) Ensure that the replaced words or phrases are distinct from the original and not synonyms; (b) Maximize the differences between the revised and original sentences by replacing as many words as possible; (c) Prioritize the replacement of words that do not have synonyms, as it is challenging to ensure that all synonyms are correctly replaced for words with multiple synonyms; (d) After replacement, the resulting statement may be inconsistent with common knowledge; (d) After replacement, it’s OK that the resulting statement may be inconsistent with common knowledge, as long as the answer can be derived from the context and that all information related to this answer is consistent within the context.
Appendix C Detailed Evaluation Results
The detailed results on each dataset of the single-hop QA task type and multi-hop QA task type are shown in Figure A8 and Figure A9, respectively. We can see that (i) Among the multi-hop QA datasets, loogle-CR-mixup and loogle-MR-mixup are particularly challenging. Future research is needed to improve the ability to aggregate multiple pieces of evidence from a long context with distracting and confusing facts. (ii) For single-hop QA datasets, as expected, LLMs can achieve higher scores on datasets without CFI, including loogle-SD-mixup and cmrc-mixup. (iii) Several LLMs, namely ChatGLM3-6B-32k, BlueLM-7B-32k-Chat, and Yi-6B-200k, can achieve relatively high scores on factrecall-en. This indicates that the NIAH task might not be challenging enough, emphasizing the need to evaluate LLMs on other tasks, particularly multi-hop QA datasets. (iv) The performance gap between LLMs on factrecall-en and factrecall-zh is especially large, and some open-source LLMs with relatively small context sizes, namely Llama2-7B-Chat-hf (4k context window size), Qwen-7B-8k-Chat, and Vicuna-7B-16k-v1.5, even get near-zero scores. (v) A few LLMs have unbalanced performances on Chinese and English datasets, as illustrated by the results on multifieldqa-en-mixup and multifieldqa-zh-mixup. The detailed scores of all models on 5 length levels of all sub-datasets are shown in Table A6 A7.
Figure A8: Average scores across all length levels of 14 LLMs on single-hop QA datasets.
Figure A9: Average scores across all length levels of 14 LLMs on multi-hop QA datasets.
Dataset Len. Ch. Bl. Yi. Lo. Qw. Vi. Ll.4 Ll.8 Ll.M GPT3. GPT4. Ll.31 Qw.25 Ms.
dureader -mixup $16k$ 23.99 19.40 2.87 13.44 11.82 9.67 7.21 16.39 18.06 8.01 19.14 20.22 18.28 21.46
$32k$ 25.21 19.74 2.98 11.57 12.80 7.65 5.42 13.08 15.86 5.26 13.64 22.38 19.16 21.00
$64k$ 22.01 14.44 2.88 9.23 10.48 6.62 5.59 10.24 15.16 4.26 12.66 19.22 17.82 -
$128k$ 17.94 10.95 2.36 9.51 8.15 6.25 4.78 5.30 14.46 3.30 8.19 9.90 18.24 -
$256k$ 8.72 8.51 3.06 7.96 8.65 5.70 4.45 4.46 10.64 3.50 6.71 5.11 16.27 -
loogle-CR -mixup $16k$ 14.41 9.01 8.25 11.25 5.48 5.00 3.69 8.63 12.56 10.04 12.68 30.05 20.33 25.02
$32k$ 14.10 7.36 8.83 11.17 3.30 4.25 3.29 8.74 11.05 8.39 10.40 26.24 20.25 23.29
$64k$ 9.92 3.81 4.73 9.31 3.82 3.76 3.13 2.78 8.64 5.58 6.48 26.81 19.08 -
$128k$ 6.95 2.40 4.05 6.19 1.14 1.99 2.19 0.26 5.81 3.08 2.83 20.72 12.56 -
$256k$ 5.46 2.60 3.23 5.03 1.94 1.28 0.81 0.49 4.54 3.37 3.91 16.61 7.31 -
loogle-MR -mixup $16k$ 15.83 4.90 6.94 10.53 4.93 5.17 3.37 10.39 13.73 12.95 12.24 23.60 19.14 20.62
$32k$ 11.62 3.14 7.67 9.51 2.95 3.83 2.20 7.14 10.9 7.03 7.83 22.54 16.12 19.63
$64k$ 7.00 1.68 2.69 3.04 2.37 0.96 2.05 3.89 7.82 6.23 6.26 21.57 16.52 -
$128k$ 7.24 2.46 3.44 4.05 1.80 0.55 1.04 2.37 5.93 2.13 2.30 18.75 12.06 -
$256k$ 3.82 2.19 1.32 3.01 1.46 1.06 0.33 0.4 4.63 1.00 0.90 15.83 7.74 -
hotpotwikiqa -mixup $16k$ 16.98 19.31 23.55 11.57 2.78 2.63 3.99 12.14 17.67 11.96 13.51 27.53 31.69 30.08
$32k$ 14.76 14.07 18.94 10.71 1.89 2.19 1.30 7.37 17.17 6.66 10.62 30.86 31.96 28.93
$64k$ 9.02 9.63 9.94 4.77 2.27 2.05 1.84 2.34 13.37 3.27 6.67 27.57 26.80 -
$128k$ 8.31 7.71 7.66 5.49 2.37 1.04 0.81 3.86 15.02 4.23 4.13 17.97 21.46 -
$256k$ 6.68 5.40 2.01 2.37 1.82 1.85 0.75 2.17 10.88 3.30 2.36 8.07 14.07 -
Table A6: Overall results of multi-hop QA tasks in LV-Eval. The abbreviations “Ch.”, “Bl.”, “Yi.”, “Lo.”, “Qw.”, “Vi.”, “Ll.4”, “Ll.8”, “Ll.M”, “Ll.31”, “Qw.25”, “Ms.”, “GPT3.”, and “GPT4.” stand for ChatGLM3-6B-32k, BlueLM-7B-32k-Chat, Yi-6B-200k, LongChat-7B-32k-v1.5, Qwen-7B-8k-Chat, Vicuna-7B-16k-v1.5, Llama2-7B-Chat-hf, Llama3-8B-Instruct, Llama3-8B-1M, Meta-Llama-3.1-70B-Instruct, Qwen2.5-72B-Instruct-128k, moonshot-v1-128k, GPT-3.5-16k, and GPT-4-8k, respectively.
Dataset Len. Ch. Bl. Yi. Lo. Qw. Vi. Ll.4 Ll.8 Ll.M GPT3. GPT4. Ll.31 Qw.25 Ms.
lic-mixup $16k$ 24.15 20.75 5.37 15.45 6.05 8.34 2.48 9.16 15.27 7.65 13.69 15.46 30.96 35.57
$32k$ 22.27 12.68 6.25 10.02 6.07 4.81 0.99 6.57 15.62 4.42 5.86 17.02 28.14 34.46
$64k$ 14.33 5.00 7.19 4.54 4.21 2.52 0.48 1.80 14.85 3.07 3.23 12.77 27.65 -
$128k$ 8.30 3.03 5.56 2.47 4.34 2.36 0.42 2.52 11.86 0.87 1.90 2.17 27.71 -
$256k$ 6.07 4.11 6.24 2.14 3.19 1.99 0.73 1.30 7.96 1.65 1.70 1.32 15.20 -
loogle-SD -mixup $16k$ 41.82 34.34 39.56 27.42 10.54 8.79 6.75 25.08 39.53 31.67 27.01 63.25 58.13 59.83
$32k$ 30.31 15.10 36.48 18.21 4.70 4.90 2.61 12.56 31.45 18.56 14.01 61.95 57.02 57.85
$64k$ 19.07 4.95 31.71 12.09 2.40 3.07 2.58 7.34 28.45 10.41 8.00 53.97 56.38 -
$128k$ 11.34 5.32 25.71 9.11 3.25 4.24 2.04 4.85 18.81 5.74 5.14 35.63 41.99 -
$256k$ 8.92 5.41 12.37 5.97 3.02 2.39 1.24 0.91 8.37 3.56 1.48 22.45 23.95 -
cmrc-mixup $16k$ 51.21 45.89 1.05 20.99 11.13 11.75 3.85 15.16 20.25 12.19 14.67 53.95 35.63 46.42
$32k$ 46.34 19.53 0.35 10.77 5.32 6.55 1.08 6.77 19.83 6.00 3.33 49.82 33.01 46.80
$64k$ 20.71 10.66 0.84 8.97 4.68 5.04 1.72 4.82 17.27 3.57 5.31 50.72 35.38 -
$128k$ 14.16 7.06 1.58 3.77 3.81 2.75 1.64 1.78 13.46 2.73 3.81 5.32 35.42 -
$256k$ 8.38 4.51 2.54 3.75 4.09 4.13 1.54 1.73 5.66 1.32 2.68 2.94 22.38 -
multifieldqa -en-mixup $16k$ 25.40 11.82 10.01 12.02 7.66 6.29 8.81 16.33 21.30 18.78 19.00 39.46 29.90 33.44
$32k$ 12.78 6.34 9.24 7.58 3.61 4.32 5.55 9.60 17.05 11.59 12.69 38.61 27.96 34.19
$64k$ 12.32 8.38 8.83 7.84 5.23 2.79 1.58 6.15 18.68 7.38 8.30 32.19 26.87 -
$128k$ 9.89 5.29 5.98 3.11 3.64 2.51 2.54 6.63 17.27 7.95 7.25 26.06 21.21 -
$256k$ 4.24 4.78 4.69 4.22 2.44 1.28 1.49 3.20 14.42 3.21 3.54 19.05 17.10 -
multifieldqa -zh-mixup $16k$ 32.38 22.05 2.85 9.81 8.82 5.82 4.72 18.73 21.69 18.94 17.61 39.16 40.65 39.05
$32k$ 24.48 17.64 0.75 8.82 5.68 4.45 1.21 13.60 13.46 12.21 11.18 32.80 36.28 31.71
$64k$ 20.97 7.36 1.89 3.23 3.01 2.03 0.68 6.13 11.31 6.29 4.99 30.77 31.79 -
$128k$ 10.08 5.90 2.11 3.54 2.84 0.88 0.24 1.52 9.28 2.94 1.76 24.42 30.71 -
$256k$ 7.05 4.48 1.58 3.92 2.52 1.26 0.56 2.62 7.79 2.15 0.92 19.37 25.95 -
factrecall-en $16k$ 91.50 58.5 24.88 9.22 1.77 0 1.08 2.72 68.00 8.25 23.4 85.20 98.50 81.38
$32k$ 89.00 32.17 23.09 14.33 1.12 0 0.46 2.03 67.17 3.27 11.84 88.36 96.00 86.17
$64k$ 46.00 15.50 24.96 8.31 0.71 0 0.31 0.61 73.00 1.80 5.21 93.33 94.00 -
$128k$ 24.00 9.00 22.04 7.86 0.18 0.25 0.23 0.15 78.83 0.60 4.03 81.88 88.00 -
$256k$ 12.50 5.00 16.44 6.00 0.22 0.20 0.15 0 58.00 0.45 1.79 47.96 52.50 -
factrecall-zh $16k$ 0 19.00 25.73 7.20 15.75 0 0 2.18 0 14.51 28.03 17.00 83.00 67.19
$32k$ 2.00 37.00 16.86 5.00 6.00 0 0 2.03 0.14 6.70 15.24 17.50 77.50 71.14
$64k$ 12.50 20.00 12.41 3.50 3.50 0 0 1.09 0 2.49 8.08 6.00 77.00 -
$128k$ 9.00 12.50 10.13 3.70 1.50 0 0 0.32 0 1.72 3.58 5.53 76.50 -
$256k$ 7.00 5.50 4.62 2.00 0.50 0 0 0.21 0 0.98 2.00 10.44 47.06 -
Table A7: Overall results of single-hop QA tasks in LV-Eval. The abbreviations “Ch.”, “Bl.”, “Yi.”, “Lo.”, “Ll.32”, “Qw.”, “Vi.”, “Ll.4”, “Ll.8”, “Ll.M”, “Ll.31”, “Qw.25”, “Ms.”, “GPT3.”, and “GPT4.” stand for ChatGLM3-6B-32k, BlueLM-7B-32k-Chat, Yi-6B-200k, LongChat-7B-32k-v1.5, Qwen-7B-8k-Chat, Vicuna-7B-16k-v1.5, Llama2-7B-Chat-hf, Llama3-8B-Instruct, Llama3-8B-1M, Meta-Llama-3.1-70B-Instruct, Qwen2.5-72B-Instruct-128k, moonshot-v1-128k, GPT-3.5-16k, and GPT-4-8k, respectively.
Appendix D Detailed Ablation Results
The detailed ablation results of CFI and KPR are shown in Table A8 A9 A10.
@@ Appendix D Detailed Ablation Results
Table A8: Ablation results on hotpotwikiqa-mixup for confusing facts insertion (CFI) and keyword and phrase replacement (KPR).
@@
Table A9: Ablation results on multifieldqa-en-mixup for confusing facts insertion (CFI) and keyword and phrase replacement (KPR).
@@
Table A10: Ablation results on multifieldqa-zh-mixup for confusing facts insertion (CFI) and keyword and phrase replacement (KPR).
@@
Table A11: Ablation results on factrecall-en and factrecall-zh for confusing facts insertion (CFI) and keyword and phrase replacement (KPR).
Appendix E Samples in LV-Eval
For completeness of our manuscript, we show some data samples of factrecall-en and factrecall-zh in Figure A10 A11.
@@ Appendix E Samples in LV-Eval
Appendix F Examples of Failure Cases
Figure A12 shows some failure cases of Llama-3-8b-Instruct when confusing facts exist in factrecall-en and factrecall-zh. In factrecall-zh-16k, all of Llama-3-8b-Instruct’s responses were misled by the confusing fact, that is “David Beckham”, whereas in the factrecall-en-16k dataset, only 32% of the responses were misled by the confusing fact. This suggests that the model’s anti-interference ability may vary significantly across languages.
Figure A13 shows a failure case of multi-hop reasoning in a QA task. In the 65th test sample of the hotpotwikiqa-mixup-16k dataset, the question asks, What is the date of death of the director of the film Nallavan Vazhvan?. Answering this question requires multi-hop reasoning: The model needs to first extract the director’s name from ### Passage 30, which discusses the film Nallavan Vazhvan, and then retrieve the final answer from ### Passage 15 which provides biographical details. However, the model incorrectly pulls the answer from ### Passage 27, which describes another director, giving a vague statement that the director was involved in the production of over 60 films. The model mistakenly interprets this as relevant information for answering the question. Additionally, the incorrect response may have been influenced by the fact that the director’s first name in the film’s introduction was abbreviated, preventing the model from retrieving the correct answer via exact matching. This example indicates that even a powerful model like Llama 3 struggles to accurately understanding entity relationships in long-context, multi-step reasoning tasks and is easily misled by superficially relevant but ambiguous information.
Figure A12: An example of model’s failure cases caused by confusing facts in factrecall-en and factrecall-zh. A list of model responses is shown in the figure.
Figure A13: A failure case of multi-hop reasoning in hotpotwikiqa-mixup-16k.